TIFFDITHER(1) TIFFDITHER(1) NAME tiffdither−convertagreyscaleimagetobilevelusingdithering SYNOPSIS tiffdither[options]input.tifoutput.tif DESCRIPTION tiffditherconvertsasinglechannel8-bitgreyscaleimagetoabilevelimageusingFloyd-Steinberg errorpropagationwiththreholding. OPTIONS −cSpecifythecompressiontousefordatawrittentotheoutputfile:nonefornocompression,packbitsforPackBitscompression,lzwforLempel-Ziv&Welchcompression,zip forDeflatecompression,g3forCCITTGroup3(T.4)compression,andg4forCCITT Group4(T.6)compression.Bydefaulttiffditherwillcompressdataaccordingtothe valueoftheCompressiontagfoundinthesourcefile. TheCCITTGroup3andGroup4compressionalgorithmscanonlybeusedwithbilevel data. Group3compressioncanbespecifiedtogetherwithseveralT.4-specificoptions:1dfor 1-dimensionalencoding,2dfor2-dimensionalencoding,andfilltoforceeachencoded scanlinetobezero-filledsothattheterminatingEOLcodeliesonabyteboundary. Group3-specificoptionsarespecifiedbyappendinga‘‘:’’-separatedlisttothe‘‘g3’’option; e.g.−cg3:2d:filltoget2D-encodeddatawithbyte-alignedEOLcodes. LZWcompressioncanbespecifiedtogetherwithapredictorvalue.Apredictorvalueof2 causeseachscanlineoftheoutputimagetoundergohorizontaldifferencingbeforeitis encoded;avalueof1forceseachscanlinetobeencodedwithoutdifferencing.LZW-spe-cificoptionsarespecifiedbyappendinga‘‘:’’-separatedlisttothe‘‘lzw’’option;e.g.−c lzw:2forLZWcompressionwithhorizontaldifferencing. −fSpecifythebitfillordertouseinwritingoutputdata.Bydefault,tiffditherwillcreatea newfilewiththesamefillorderastheoriginal.Specifying−flsb2msbwillforcedatato bewrittenwiththeFillOrdertagsettoLSB2MSB,while−fmsb2lsbwillforcedatato bewrittenwiththeFillOrdertagsettoMSB2LSB. −tSetthethresholdvaluefordithering.Bydefaultthethresholdvalueis128. NOTES Theditheralgorithmistakenfromthetiffmedian(1)program(writtenbyPaulHeckbert). SEEALSO pal2rgb(1),fax2tiff(1),tiffinfo(1),tiffcp(1),tiff2bw(1),libtiff(3) October15,1995 1